Project Management and Quality Control Question

Question 2)

A gardening supply company manufactures two types of mulch bins – the Compacto and the Enviro. The production of each bin involves a certain amount of moulding, pressing and assembly as indicated below.

Department

Compacto

Enviro

Moulding

5 hours

6 hours

Pressing

8 hours

5 hours

Assembly

9 hours

14 hours

Next month, the company has maximum available capacities of 630 hours for assembly, 300 hours for moulding and 400 hours for pressing. The company has advance orders for 15 Compacto mulch bins and for 10 Enviro mulch bins and is committed to supplying these next month. The same profit of $6,000 is made from the sale of each mulch bin.

(i) How many Compacto and Enviro mulch bins should be manufactured next month to fulfill advance orders and maximize profits? How much profit will be made? (Assume that partially completed mulch bins can be manufactured but that only completed bins can be sold for profit. Bins on advance order have not yet been paid for.) Solve both graphically and mathematically.

(ii) At the optimum production mix for next month, which if any, of the three departments has any excess capacity, i.e. hours available that are not used? Indicate the amount of any excess capacity in hours
Answer #1

(i) Solution using graphical method is following:

Number of Compacto mulch bins to be manufactured (X1) = 28.89

Number of Enviro mulch bins to be manufactured (X2) = 10

Total profit = 28*6000+10*6000 = $ 228,000

(ii)

Excess capacity in Moulding dept = 630 - (5*28.89 + 6*10) = 425.55 hours

Excess capacity in Pressing dept = 300 - (8*28.89 + 5*10) = 18.88 hours

Excess capacity in Assembly dept = 400 - (9*28.89 + 14*10) = 0 hours